OpenStudy (anonymous):

x^3−4x^2+x−6 divided by (x+2). If (x+2) is a factor, determine the remaining factors.

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

your attempt?

OpenStudy (anonymous):

idk what the second part means

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

do you know long division? or synthetic?

OpenStudy (anonymous):

long'

OpenStudy (anonymous):

I have the anwer

OpenStudy (loser66):

show me your work

OpenStudy (anonymous):

Its to much to type

OpenStudy (anonymous):

the answer I got is x^2-2x-3 @Loser66

OpenStudy (anonymous):

remainder 12

OpenStudy (loser66):

hey, how can I know whether it is your work or you just copy from the back of the book? show your work, please, use draw box

OpenStudy (anonymous):

trust me, I dont have the answers its a worksheet

OpenStudy (loser66):

hahaha, ok, let me check.

OpenStudy (anonymous):

or else I wouldnt be asking:P

OpenStudy (loser66):

hhhmmm how to say. :) I don't get what you got. let work together

OpenStudy (anonymous):

I dont get the question

OpenStudy (loser66):

cannot use the draw box ,no choice for us, give me few minutes, I work on the paper, scan and post it . OK

OpenStudy (anonymous):

ok

OpenStudy (loser66):

can you see it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

dont submit a pdf

OpenStudy (anonymous):

I can't open pdfs

OpenStudy (loser66):

One more try x^2 -6x +13 |------------------ x+2 | x^3 - 4x^2 +x -6 - x^3 +2x^2 ---------------- 0 -6x^2 +x - -6x^2 -12x --------------------- 0 -13x -6 - -13x +26 ----------------- 0 -32 so the remainder is -32

OpenStudy (loser66):

see it?

OpenStudy (loser66):

hey, I did all my best, give me your cupcake and I log out. hehehhe. joke. Do you get it?

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

YEAH. means you got it, right?

OpenStudy (anonymous):

but what is meant by the second part of the question?

OpenStudy (loser66):

it means the remainder factor is \(\dfrac{-32}{x+2}\)

OpenStudy (anonymous):

huuuh?

OpenStudy (loser66):

I know it didn't show, -32 in the numerator, (x+2) in the denominator. Is it clear?

OpenStudy (loser66):

that is the remainder factor. got me? write it out on the paper

OpenStudy (anonymous):

Help me with this one 3x^3−5x^2−16x+12 divided by (x+2). If (x+2) is a factor, determine the remaining factors.

OpenStudy (loser66):

hey, hey, it's exactly what we 've just done. DO IT. show me you got something from my guiding. I can check. I don't want to be your free answer service. you do it by yourself now. Show me your brain, please

OpenStudy (anonymous):

k wait

OpenStudy (anonymous):

I got .. 3x^2 -11x +6 Reminder 0

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

?????

OpenStudy (loser66):

nope, redo. If you want me trust you, show the whole work like what I did. I could do it, why couldn't you? Moreover, it's your problem, right? One more try, please

OpenStudy (anonymous):

its wrong?

OpenStudy (loser66):

IT'S NOT CORRECT

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

oh, yea, you are right , I am sorry. (wonder if I don't say it is not correct, is there anyway else to force you show the work?? ) hehehe.

OpenStudy (anonymous):

but the second part of the question???

OpenStudy (loser66):

it's me, I am a rude person. don't you know it? hehehe , the second part is "no remainder fraction"

OpenStudy (anonymous):

Complete the square to find the vertex of each quadratic function. State the vertex. y=2x^2+8x−2

OpenStudy (loser66):

your idea?

OpenStudy (anonymous):

x1=−2−5√ x2=−2+5√

OpenStudy (anonymous):

oops square root 5

OpenStudy (loser66):

nope, not that way

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

help me

OpenStudy (loser66):

I am thinking the easiest way to show you

OpenStudy (loser66):

question : 2x^2 or (2x)^2

OpenStudy (anonymous):

y=2x^2+8x−2

OpenStudy (loser66):

we got trouble then. hehehe. let see what I can do for you

OpenStudy (loser66):

ok, factor 2 out, you have y = 2 ( x^2+4x -1)

OpenStudy (loser66):

follow me carefully and stop me when you don't get me. Ok? i will post step by step

OpenStudy (anonymous):

ok

OpenStudy (loser66):

I just pick the ( ) part to solve first: x^2 + 4x + 4 -4 -1 = (x +2)^2 -5 got me ??

OpenStudy (loser66):

no?

OpenStudy (anonymous):

kind of

OpenStudy (loser66):

question?

OpenStudy (anonymous):

are you doing perfect squares?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

completeing *

OpenStudy (anonymous):

ok

OpenStudy (loser66):

got it? right?

OpenStudy (anonymous):

ya

OpenStudy (loser66):

not finish yet, just first step

OpenStudy (anonymous):

ok

OpenStudy (loser66):

now, put it back to y y = 2 ( ) , right? and from the previous post , I have ( ) = (x+2) ^2 -5 so y = 2 ((x+2)^2 -5) = 2 (x+2)^2 -10 got it?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

the vertex?

OpenStudy (loser66):

so, we have the form of vertex form y = a (x -h)^2 + k where h,k is vertex of parabola your is y = 2 (x +2) ^2 -10 , can you tell me h =?? k =??

OpenStudy (anonymous):

-2,10

OpenStudy (loser66):

-2, -10, not 10

OpenStudy (loser66):

Yes or No?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

vertex is -2,-10

OpenStudy (anonymous):

righttt????

OpenStudy (loser66):

YES hehehe.... this rude guy is acceptable, right? hehehe

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

y=-x^2+5x−3

OpenStudy (anonymous):

another one

OpenStudy (loser66):

I don't know how to solve it, friend, hehehe... need rest.

OpenStudy (anonymous):

one more!!!

OpenStudy (loser66):

post in new one, I will tag Mertsj